# Higher-Order Networks Representation and Learning: A Survey

### 29 Feb 2024 v1

@@ 1 Introduction
Organization. The rest of this paper is structured as follows. Section 2 introduces necessary dyadic graph basics. Sections 3, 4, and 5 introduce foundations, algorithms, and applications of network motifs, simplicial complexes, and hypergraphs, respectively. Section 6 summarizes existing datasets and tools that have been used in higher-order network studies. We present challenges and future direction and conclude in Section 7.
@@ 5.1.2 Tensor Representation
CP (CANDECOMP/PARAFAC) Decomposition (71; 72) is also called tensor rank decomposition or Canonical Polyadic Decomposition (CPD). The CP decomposition factorizes a tensor into a sum of component vectors. For example, a tensor $\mathbf{T}\in\mathbb{R}^{I\times J\times K}$ can be decomposed as
@@ 5.2.2 Generative Models
Furthermore, a later study extends the degree-corrected stochastic blockmodel (85), which is a generative model of graphs with both community structure and degree sequences, to hypergraphs (86). For hyperedge candidates, the authors introduce an affinity function to compute the wiring possibility based on the group memberships of their nodes. Basically, more nodes in the same group have a higher probability of forming hyperedges. Three estimates —the affinity function, node labels, and node degrees, are alternatively learned by optimizing a likelihood function. To solve this objective, the authors propose an ‘All-or-Nothing’ (AON) Louvain-type algorithm under the assumption that hyperedges are expected to lie fully within the cluster. Experimental results on both synthetic and empirical data validate the efficiency and accuracy of the framework.
@@
Package Name Environment Description Official Link Simplicial Python topology, homology, filtrations https://simplicial.readthedocs.io/en/latest/ Javaplex (96) Matlab/Java persistent homology, filtrations https://github.com/appliedtopology/javaplex Ripser (97) C++ persistent homology, Vietoris–Rips filtrations https://github.com/Ripser/ripser simplextree R topology https://github.com/peekxc/simplextree Simplicial.jl Julia simplicial complexes, directed complexes https://github.com/nebneuron/Simplicial.jl simplicial-complex JavaScript structural and topological operations https://www.npmjs.com/package/simplicial-complex Dionysus 2 C++ persistent homology https://mrzv.org/software/dionysus2/ DIPHA C++ distributed, persistent homology https://github.com/DIPHA/dipha Perseus (98) C++ persistent homology https://people.maths.ox.ac.uk/nanda/perseus/ Moise Maple homology groups https://www.math.drexel.edu/~ahicks/Moise/
@@ 5.2.5 Hypergraph Neural Networks
Graph neural networks have been generalized beyond pairwise interactions in modeled as hypergraphs. Hypergraphs can be applied to either (1) model higher-order graph data as input matrices, such as the incident matrix of hypergraphs; or to (2)build multilayer neural network structures but using higher-order forward- and back-propagation instead. Many state-of-the-art models, especially for graph neural networks, are extended to hypergraphs. Examples include hyper- models such as HGNN (106), HGAT (107; 108), MHCN (109), and HGCN (110; 111). For a comprehensive survey on graph neural networks, readers are referred to the survey by Thomas et al. (112).
@@
Network Motifs Simplicial Complexes Hypergraphs Statistical Significance (11)(17)(15)(16)(41) Graph Classification (12)(13) Network Modeling (42)(18) (50)(63)(65)(66) (82)(83)(86) Clustering (38)(43)(40)(38)(20) (68) (74)(99)(100)(101)(102) Representation Learning (44)(45) (120)(121) Link Prediction (46) (20) (105)(120) Persistent Homology (52)(53)(54)(55)(56)(57)(58)(59)(60)(61)(62) Analysis Tools and Measurements (40)(38) (49) (78)(79)(80)(81) Recommender System (103)(104) Neural Networks (106)(107)(108)(109)(110)(111)
@@ 6.2 Higher-Order Datasets
We provided dataset resources with higher-order interactions:
  • ARB Data [^2] https://www.cs.cornell.edu/~arb/data/: A dataset repository collected by Austin R. Benson. Most are higher-order networks from various fields, including temporal hypergraphs, labeled hypergraphs, etc.
  • LINQS [^3] https://linqs.soe.ucsc.edu/: A collection of several relational datasets. Many of them are collaboration networks, which naturally include higher-order interactions.
  • Twitter Data [^4] https://data.world/datasets/twitter: Tweets can be modeled as higher-order networks by taking hashtags as nodes and co-appearances as edges. Besides these, one can search twitter datasets online for any specific interest or collect data using APIs.
  • Temporal Co-authorship  [^5] https://github.com/kswoo97/pcl: Three large-scale hypergraph datasets in both static and temporal forms (122).
@@ 7.2 Machine Learning Applications
Representation Learning: Representation learning is a powerful tool for transforming high-dimensional data into fixed-size vectors and has been extremely successful for downstream machine learning tasks, such as node classification, link prediction, among others. However, not many representation learning methods are introduced for higher-order networks. Hence, there is a significant demand for representation learning methods that can embed higher-order graphs both at the node-level (120; 121) and the graph-level.

### 09 Apr 2024 v2

@@ 1 Introduction
Scope and Organization. Compared to other surveys, our goal is to provide a succinct yet broad and comprehensive survey that focuses on higher-order networks. Abstract network topologies are categorized on the basis of the data type and applications. The surveys aims to assist researchers in identifying the appropriate methods, resources, and higher-order tools for their research tasks.
The rest of this paper is structured as follows. Section 2 introduces necessary dyadic graph basics. Sections 3, 4, and 5 introduce foundations, algorithms, and applications of network motifs, simplicial complexes, and hypergraphs, respectively. Section 6 summarizes existing datasets and tools that have been used in higher-order network studies. We present challenges and future direction and conclude in Section 7.
@@ 4.2.5 Network Analysis Tools
Advanced network representation learning methods have also been extended to the simplicial complex. Hajij et al. (69) use the autoencoder to perform simplex-level embedding. The encode function ($X\rightarrow\mathbb{R}^{d}$) maps each simplex to a fixed vector. The decode function ($\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$) maps each pair of simplices to a similarity score that reflects the relationship between the two simplices. An example of a user-defined similarity is the simplex-level adjacency matrix. Finally, the representation of the whole simplicial complex is obtained by a weighted sum of the simplex-level representations.
@@ 5.1.2 Tensor Representation
CP (CANDECOMP/PARAFAC) Decomposition (72; 73) is also called tensor rank decomposition or Canonical Polyadic Decomposition (CPD). The CP decomposition factorizes a tensor into a sum of component vectors. For example, a tensor $\mathbf{T}\in\mathbb{R}^{I\times J\times K}$ can be decomposed as
@@ 5.2.2 Generative Models
Furthermore, a later study extends the degree-corrected stochastic blockmodel (86), which is a generative model of graphs with both community structure and degree sequences, to hypergraphs (87). For hyperedge candidates, the authors introduce an affinity function to compute the wiring possibility based on the group memberships of their nodes. Basically, more nodes in the same group have a higher probability of forming hyperedges. Three estimates —the affinity function, node labels, and node degrees, are alternatively learned by optimizing a likelihood function. To solve this objective, the authors propose an ‘All-or-Nothing’ (AON) Louvain-type algorithm (88) under the assumption that hyperedges are expected to lie fully within the cluster. Experimental results on both synthetic and empirical data validate the efficiency and accuracy of the framework.
@@
Package Name Environment Description Official Link Simplicial Python topology, homology, filtrations https://simplicial.readthedocs.io/en/latest/ Javaplex (98) Matlab/Java persistent homology, filtrations https://github.com/appliedtopology/javaplex Ripser (99) C++ persistent homology, Vietoris–Rips filtrations https://github.com/Ripser/ripser simplextree R topology https://github.com/peekxc/simplextree Simplicial.jl Julia simplicial complexes, directed complexes https://github.com/nebneuron/Simplicial.jl simplicial-complex JavaScript structural and topological operations https://www.npmjs.com/package/simplicial-complex Dionysus 2 C++ persistent homology https://mrzv.org/software/dionysus2/ DIPHA C++ distributed, persistent homology https://github.com/DIPHA/dipha Perseus (100) C++ persistent homology https://people.maths.ox.ac.uk/nanda/perseus/ Moise Maple homology groups https://www.math.drexel.edu/~ahicks/Moise/ TopoEmbedX (101) Python representation learning https://github.com/pyt-team/TopoEmbedX
@@ 5.2.5 Hypergraph Neural Networks
Graph neural networks have been generalized beyond pairwise interactions in modeled as hypergraphs. Hypergraphs can be applied to either (1) model higher-order graph data as input matrices, such as the incident matrix of hypergraphs; or to (2)build multilayer neural network structures but using higher-order forward- and back-propagation instead. Many state-of-the-art models, especially for graph neural networks, are extended to hypergraphs. Examples include hyper- models such as HGNN (109), HGAT (110; 111), MHCN (112), and HGCN (113; 114). For a comprehensive survey on graph neural networks, readers are referred to the survey by Thomas et al. (115).
@@
Network Motifs Simplicial Complexes Hypergraphs Statistical Significance (11)(17)(15)(16)(41) Graph Classification (12)(13) Network Modeling (42)(18) (50)(63)(65)(66) (83)(84)(87) Clustering (38)(43)(40)(38)(20) (68) (75)(102)(103)(104)(105) Representation Learning (44)(45) (69) (123)(124) Link Prediction (46) (20) (108)(123) Persistent Homology (52)(53)(54)(55)(56)(57)(58)(59)(60)(61)(62) Analysis Tools and Measurements (40)(38) (49) (79)(80)(81)(82) Recommender System (106)(107) Neural Networks (109)(110)(111)(112)(113)(114)
@@ 6.2 Higher-Order Datasets
We summarize some dataset resources with higher-order interactions:
  • ARB Data [^2] https://www.cs.cornell.edu/~arb/data/: A dataset repository (19 datasets–4 with millions of nodes, 10 with thousands of nodes, and 5 with hundreds of nodes) collected by Austin R. Benson. Most are higher-order networks from various fields, including temporal and labeled hypergraphs.
  • LINQS [^3] https://linqs.soe.ucsc.edu/: A collection of 11 relational datasets (1 with a million nodes, others are thousands of nodes or less). Many of them are collaboration networks, which naturally include higher-order interactions.
  • Twitter Data [^4] https://data.world/datasets/twitter: Tweets can be modeled as higher-order networks by taking hashtags as nodes and co-appearances as edges (require preprocessing). Besides these, one can search twitter datasets online for any specific interest or collect data using APIs.
  • Temporal Co-authorship  [^5] https://github.com/kswoo97/pcl: Three large-scale (sizes: 27 million / 13 million / 41 thousand nodes) hypergraph datasets in both static and temporal forms (125).
6.3 Expected Time Complexities
While time complexity of exploring higher-order networks can vary across topologies and algorithms, some time complexities are typically expected for some basic higher-order algorithms. Here, we briefly list some expected time complexities for analyzing higher-order networks.
Motif counting: The time complexity of counting motifs depends highly on the structural complexity of given motifs as the essential algorithm for finding motifs involves checking for subgraph isomorphism, which is known to be NP-complete. Most fast motif-finding algorithms focus on size 3 or 4 motifs. For example, counting motifs of size 3 (triangles) can be solved in $O(|E|d_{max})$ (126) and counting motifs of size 4 can be solved in $O(|E|d_{max}+|E|^{2})$ (127), where $d_{max}$ is the maximum degree in the graph.
Homology groups: The time complexity of computing homology groups of the simiplical complex is $O(n^{\omega})$, where $n$ is the number of simplices and the exponent $\omega\leq 2.4$ (128). Such an acceptable and stable complexity facilitates the wide usage of homology methods.
@@ 7.1 Data Source and Modeling
Interpretability and Causality: Most higher-order networks studies develop algorithms and applications that explore collected higher-order data. Current research rarely aims to interpret findings in higher-order networks or understand why some high-order interactions or patterns exist. As mentioned, some patterns might reflect important real-world information beyond network structures, such as the small functional unit in brain networks (15). Interpretability is especially crucial as more black-box techniques (e.g., deep neural networks) or embedding methods are designed for higher-order networks.
@@ 7.2 Machine Learning Applications
Representation Learning: Representation learning is a powerful tool for transforming high-dimensional data into fixed-size vectors and has been extremely successful for downstream machine learning tasks, such as node classification, link prediction, among others. However, not many representation learning methods are introduced for higher-order networks. Hence, there is a significant demand for representation learning methods that can embed higher-order graphs both at the node-level (123; 124) and the graph-level.